# Directional testing for one-way MANOVA in divergent dimensions

### v2

@@ 4.2 Heteroscedastic one-way MANOVA
The performance of the directional test for hypothesis (9) in the moderate dimensional multivariate normal framework is here evaluated via Monte Carlo simulations based on 10,000 replications. The directional test is again compared with the chi-square approximations for the likelihood ratio test and its modifications proposed by Skovgaard (2001). When $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$ by Nel and Merwe (1986). The different testing approaches are evaluated in terms of empirical size. The result for $g=5,30$ are reported in Section S2.3 of the Supplemental material , and are in line with the ones available discussed below.
@@
We generate the data matrix $y_{i}$ as $n_{i}$ independent replications from a multivariate normal distribution $N_{p}(\mu_{i},\Lambda_{i}^{-1}),i=1,\dots,g$. Under the null hypothesis $H_{\psi}$, we set $\mu_{i}=0_{p}$ and use an autoregressive structure for the covariance matrices, i.e. $\Lambda_{i}^{-1}=(\sigma_{jl})_{p\times p}=(\rho_{i}^{|j-l|})_{p\times p}$, with the $\rho_{i}$ chosen to an equally-spaced sequence from 0.1 to 0.9 of length $g$. In particular, when $g=2$, $\Lambda_{1}^{-1}=(0.1^{|j-l|})_{p\times p}$ and $\Lambda^{-1}_{2}=(0.9^{|j-l|})_{p\times p}$. We show results for $p=\lceil n_{i}\rceil^{\tau}$ with $\tau=j/24$, $j\in\{6,7,\dots,22\}$, $n_{i}\in\{100,500,1000\}$ and $g=2$. Note that for $n_{i}=1000$ the simulations results are based on $5000$ replications when $j\in\{21,22\}$ due to the expensive computational cost.
@@
Figure 3: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (Nel and Merwe, 1986), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to $n_{i}=100,500,1000$, respectively ($g=3$).
@@
Figure 3 reports the empirical size at the nominal level $\alpha=0.05$ under the null hypothesis for various statistics. The directional test is always more reliable than its competitors in terms of the empirical size, even if in this case it is not exact. Skovgaard’s modifications are not as accurate when $p$ is large. Moreover, with $n_{i}$ increasing, the likelihood ratio test and the Behrens-Fisher test will not be valid and break down fast when $p=\lceil n_{i}\rceil^{\tau}$, even for moderate large values of $\tau$. Additional simulation results are showed in Table 2 for different structures of the covariance matrices: (I) identity matrix, $\Lambda_{i}^{-1}=I_{p}$; (II) compound symmetric matrix, $\Lambda^{-1}_{i}=(1-\rho_{i})I_{p}+\rho_{i}1_{p}1_{p}^{\top}$, with the same values of $\rho_{i}$ as in the autoregressive structure. Table 2 reports the empirical size at the nominal level $\alpha=0.05$ for $n_{i}=100$ and $p=\lceil n_{i}\rceil^{\tau}$, with covariance structures (I) and (II) and $g=2,30$. As expected, the likelihood ratio test performs in general very poorly, especially for larger values of $g$. Skovgaard’s modifications are much more accurate but not as much as the directional test, whose excellent performance is confirmed. When $g=2$, the approximate solution for the Behrens-Fisher test seems reliable in the identity covariance matrix case, but not under compound symmetry.
@@
Table 2: Empirical size of the directional test (DT), likelihood ratio test (LRT), two Skovgaard’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) (only for $g=2$) for hypothesis (9) at nominal level $\alpha=0.05$, with $n_{i}=100$ and $p=n_{i}^{\tau}$.
@@
4.3 Robustness to misspecification
In general, all the approaches examined so far rely on the assumed normal model and are not guaranteed to be robust under model misspecification. We can assess numerically the robustness of the various competitors using simulations. In this section, we consider three different distributions for the true generating process: multivariate $t$, multivariate skew-normal (Azzalini and Capitanio, 1999) or multivariate Laplace. More in detail, a multivariate $t$ distribution with location $0_{p}$, scale matrix $I_{p}$ and degrees of freedom $5$, a multivariate skew-normal distribution with location $1_{p}$, scale matrix $\Omega=(\omega_{jl})=(0.2)^{|j-l|}$ and shape parameter $1_{p}$, and a multivariate Laplace distribution with mean vector $1_{p}$ and identity covariance matrix. Simulation results are based on $10,000$ replications.
For hypothesis (8), Figures S5S6 show the empirical size at the nominal level $\alpha=0.05$ if the underling distribution is misspecified. We see that the directional test still maintains the hightest accuracy. For hypothesis (9), Figures 67 show the same relative pattern as that observed under the correct model specification in Figure 3. Hence, all solutions exhibit a similar behavior and seem robust to misspecification of the normal model, at least with respect to the three true data generating process we considered.
Figure 4: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (8), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=100$ and $g=3$.
Figure 5: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two Skovgaard’s modifications (Sko1 and Sko2, respectively) for hypothesis (8), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=3$.
Figure 6: Empirical size for the directional test (DT), Behrens-Fisher test (BF), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions if the true generating process, respectively, with $n_{i}=100$ and $g=2$.
Figure 7: Empirical size for the directional test (DT), Behrens-Fisher test (BF), likelihood ratio test (LRT), and two Skovgaard’s modifications (Sko1 and Sko2) for hypothesis (9), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=2$.
@@ 5 Applications
The first dataset is taken from a study on how $p=5$ characteristics of the blood (red cell count, hemoglobin concentration, hematocrit, white cell count, and plasma ferritin concentration) vary with sport, body size, and sex of the athletes (Telford and Cunningham, 1991). Our goal is to investigate whether the blood characteristics of athletes differ between endurance-related events and power-related events. The data can be found in the function ais of the R (R Core Team, 2023) package DAAG (Maindonald and Braun, 2024). We performed several tests on this dataset to study the difference in blood parameters between athletes performing endurance-related and power-related sports. Typically, the conclusions drawn by the directional approach align with those from its best competitors. As an illustration, we report here some results involving two hypotheses: (i) testing the equality of blood characteristics between athletes competing in an endurance-related event (400-meter track event, $T_{400}$) and those in a power-related event (sprint, $T_{Sprint}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{T_{Sprint}}$; and (ii) testing the equality of blood characteristics for three groups of athletes that perform endurance-related sports ($Row$, $Swim$, and $T_{400}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}$. Here we focus exclusively on data from male athletes as each blood variable is significantly different between genders (Telford and Cunningham, 1991). We also consider the variable plasma ferritin concentration on the logarithmic scale. The sample sizes of $T_{400}$, $T_{Sprint}$, $Row$ and $Swim$ groups are respectively 18, 11, 15, 13. All the testing procedures under analysis can then be used to perform one-way MANOVA, and the corresponding $p$-values are shown in Table 3.
@@
For hypothesis (i), the $p$-values in Table 3 of the directional test, the Bartlett corrected test, the two Skovgaard’s modifications and the Behrens-Fisher test do not reject the null hypothesis in both homoscedastic and heteroscedastic cases at the significance level $\alpha=0.05$, claiming that the average blood characteristics levels for male athletes competing in $T_{400}$ and $T_{Sprint}$ events are not significantly different. Instead, the central limit theorem test and the likelihood ratio test reject the null hypothesis. Note that the central limit theorem test is not reliable with small $p$ and $n_{i}$, while the likelihood ratio test is not valid if $\kappa$ and $\tau$ are large. To assess the validity of the various testing procedures in this setting, we can also examine the empirical size of the various tests. We generate $R=10,000$ samples with sizes $n_{T_{400}}=18$ and $n_{T_{Sprint}}=11$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{T_{Sprint}}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{T_{Sprint}}\bar{y}_{T_{Sprint}})/(n_{T_{400}}+n_{T_{Sprint}})$ where $\bar{y}_{T_{400}}$ and $\bar{y}_{T_{Sprint}}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}^{-1}_{0}$, the pooled sample covariance matrix for the homoscedastic case, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$ sample covariance matrices of each group, for the heteroscedastic case. The simulation results for hypothesis (i) reported in Table 4 show that the directional test is more reliable than its competitors, having empirical size closer to the nominal level $0.05$. The central limit theorem test and likelihood ratio test suffer from large bias. These results confirm the simulation outputs in Section S1.
@@
Table 3: $p$-values of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two Skovgaard’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
Table 4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two Skovgaard’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
For hypothesis (ii), the $p$-values in Table 3 of most of the methods clearly indicate to not reject the hypothesis of equal mean blood characteristics of endurance-related events ($T_{400}$, $Row$ and $Swim$) for male athletes under heteroscedasticity at significance level $\alpha=0.05$. However, in the homoscedastic case, the likelihood ratio test is the only one to reject the hypothesis at the $0.05$ significance level. Conversely, the other statistics do not reject the null hypothesis and indicate the same mean level of blood characteristics for the three groups. Even in this case, a simulation study can be performed in order to verify the reliability of the considered methods. Under the null hypothesis, we generate $10,000$ samples with sizes $n_{T_{400}}=18$, $n_{Row}=15$ and $n_{Swim}=13$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{Row}\bar{y}_{Row}+n_{Swim}\bar{y}_{Swim})/(n_{T_{400}}+n_{Row}+n_{Swim})$ where $\bar{y}_{T_{400}}$, $\bar{y}_{Row}$ and $\bar{y}_{Swim}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}_{0}^{-1}$, the pooled sample covariance matrix under the homoscedasticity, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$, sample covariance matrices under the heteroscedasticity. Simulation results are reported in Table 4. In the homoscedastic case, the two Skovgaard’s modifications are conservative. The central limit theorem test and the likelihood test, on the other hand, have slightly inflated empirical size. Hence, the $p$-values of these four methods are not reliable in this scenario. Contrarily, the directional test and the Bartlett corrected test well control the type I error. In the heteroscedastic case, the likelihood ratio test is overly inflated, whereas the remaining procedures, including the directional test, exhibit a good performance.
The second dataset studied is taken from transcript profiles of primary breast tumors to investigate the p53 status. It is based on 251 samples predominantly containing tumor tissue that underwent sequencing to identify the p53 mutation status through RNA extraction and subsequent microarray analysis. The RNA was extracted from each sample by using Affymetrix U133 oligonucleotide microarrays, targeting over 30,000 genes (Miller et al., 2005). Here we focus on a subset of genes corresponding to a signature that is considered important in breast cancer from the research of Bonavita et al. (2020), which stores the appropriately normalized expression values of $p=22$ gene products. The samples were divided into three groups based on tumor grade. After excluding two samples due to the missing grade variable, we obtain $g=3$ groups with $n_{1}=67$, $n_{2}=128$ and $n_{3}=54$ samples for grade I, grade II and grade III, respectively. Within a high dimensional framework, we have $p/\min(n_{i})=0.407$. Homoscedastic one-way MANOVA leads to the following $p$-values: $1.565\times 10^{-7}$ for the directional test, $1.966\times 10^{-12}$ for the central limit theorem test, $2.627\times 10^{-8}$ for the likelihood ratio test, $1.915\times 10^{-7}$ for the Bartlett corrected test, and $1.828\times 10^{-7}$ and $1.877\times 10^{-7}$ for the two Skovgaard’s modifications. All methods clearly indicate to reject the null hypothesis of equal breast tumors gene expression signature in case of homoscedastic groups. As already seen, the likelihood ratio test results in the smallest $p$-value, whereas its Skovgaard’s modified versions and the Bartlett corrected test are closer to the exact directional $p$-value, consistently with our simulations. To check the reliability of the various methods in the homoscedastic case, another simulation study on the gene expression data can be conducted. The empirical size based on $10,000$ replications for the directional, central limit theorem, likelihood ratio, Bartlett corrected and two Skovgaard’s modifications tests are 0.049, 0.060, 0.087, 0.049 0.048 and 0.047, respectively. Those of central limit theorem and likelihood ratio test are a little inflated. The remaining approaches, including the directional test, appear to be accurate.
Assuming heteroscedastic groups, for hypothesis (9) the directional $p$-value is $2.861\times 10^{-5}$, the $p$-value of the likelihood ratio test is $1.928\times 10^{-7}$, and the $p$-values of the two Skovgaard’s modifications are $1.772\times 10^{-5}$ and $2.079\times 10^{-5}$. The same conclusion is hence drawn by all methods. Once again, the $p$-value of the likelihood ratio test is relatively smaller than the others, and in problems of this kind such discrepancies are particularly relevant. Even under heterescedasticity a simulation study can be performed. The empirical sizes of the directional test, the likelihood ratio test and the two Skovgaard’s modifications based on $10,000$ replications are 0.051, 0.181, 0.057 and 0.054, respectively. The directional test performs very well, as expected, while the likelihood ratio test seams still not reliable.

### v3

@@ 4.2 Heteroscedastic one-way MANOVA
The performance of the directional test for hypothesis (12) in a moderate dimensional multivariate normal framework with $p=O(n_{i}^{\tau})$, $\tau\in(0,1)$ is here evaluated via Monte Carlo simulations based on 10,000 replications. The directional test is again compared with the chi-square approximations for the likelihood ratio test and its modifications proposed by skovgaard:2001. When $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$ by nel1986. The different testing approaches are evaluated in terms of empirical size. The results for $g=5,30$ are reported in Section S2.3 of the Supplementary Material and are in line with the ones discussed below.
@@
We generate the data matrix $y_{i}$ as $n_{i}$ independent replications from a multivariate normal distributions $N_{p}(\mu_{i},\Lambda_{i}^{-1}),i\in\{1,\dots,g\}$. Under the null hypothesis $H_{\psi}$, we set $\mu_{i}=0_{p}$ and use an autoregressive structure for the covariance matrices, i.e. $\Lambda_{i}^{-1}=(\sigma_{jl})_{p\times p}=(\rho_{i}^{|j-l|})_{p\times p}$, with the $\rho_{i}$ chosen as an equally-spaced sequence from 0.1 to 0.9 of length $g$. In particular, when $g=2$, $\Lambda_{1}^{-1}=(0.1^{|j-l|})_{p\times p}$ and $\Lambda^{-1}_{2}=(0.9^{|j-l|})_{p\times p}$. We show results for $p=\lceil n_{i}\rceil^{\tau}$ with $\tau=j/24$, $j\in\{6,\dots,22\}$, $n_{i}\in\{100,500\}$ and $g=2$.
@@
Figure 3: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (12) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
@@
Figure S10 reports the empirical size at the nominal level $\alpha=0.05$ under the null hypothesis for various statistics. The directional test is always more reliable than its competitors in terms of the empirical size, even if in this framework it is not exact. skovgaard:2001’s modifications are not as accurate when $p$ is large. Moreover, the likelihood ratio test and the Behrens-Fisher test break down quickly when $p=\lceil n_{i}\rceil^{\tau}$, even for moderately large values of $\tau$. Additional simulation results are shown in Table 2 for different structures of the covariance matrices: (I) identity matrix, $\Lambda_{i}^{-1}=I_{p}$; (II) compound symmetric matrix, $\Lambda^{-1}_{i}=(1-\rho_{i})I_{p}+\rho_{i}1_{p}1_{p}^{\top}$, with the same values of $\rho_{i}$ as in the autoregressive structure. Table 2 reports the empirical size at the nominal level $\alpha=0.05$ for $n_{i}=100$ and $p=\lceil n_{i}\rceil^{\tau}$, with covariance structures (I) and (II) and $g=2,30$. As expected, the likelihood ratio test performs in general very poorly, especially for larger values of $g$. skovgaard:2001’s modifications are much more accurate but not as much as the directional test, whose excellent performance is confirmed. When $g=2$, the approximate solution for the Behrens-Fisher test seems reliable in the identity covariance matrix case, but not under compound symmetry.
@@
Table 2: Empirical size of the directional test (DT), likelihood ratio test (LRT), two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) (only for $g=2$) for hypothesis (12) at nominal level $\alpha=0.05$, with $n_{i}=100$ and $p=n_{i}^{\tau}$.
@@ 5 Applications
The dataset is taken from a study on how $p=5$ characteristics of the blood (red cell count, hemoglobin concentration, hematocrit, white cell count, and plasma ferritin concentration) vary with sport, body size, and sex of the athletes (telford1991sex). Our goal is to investigate whether the blood characteristics of athletes differ between endurance-related events and power-related events. The data can be found in the function ais of the R (Rcode) package DAAG (daag24). We performed several tests on this dataset to study the difference in blood parameters between athletes performing endurance-related and power-related sports. Typically, the conclusions drawn by the directional approach align with those from its best competitors. As an illustration, we report here some results involving two hypotheses: (i) testing the equality of blood characteristics between athletes competing in an endurance-related event (400-meter track event, $T_{400}$) and those in a power-related event (sprint, $T_{Sprint}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{T_{Sprint}}$; and (ii) testing the equality of blood characteristics for three groups of athletes that perform endurance-related sports ($Row$, SwimSwim, and $T_{400}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}$. Here we focus exclusively on data from male athletes as each blood variable is significantly different between genders (telford1991sex). We also consider the variable plasma ferritin concentration on the logarithmic scale. The sample sizes of $T_{400}$, $T_{Sprint}$, $Row$ and $Swim$ groups are respectively 18, 11, 15, 13. All the testing procedures under analysis can then be used to perform one-way MANOVA, and the corresponding $p$-values are shown in Table 3.
@@
For hypothesis (i), the $p$-values in Table 3 of the directional test (homoscedastic: 0.060; heteroscedastic: 0.097), the Bartlett corrected test (homoscedastic: 0.059), the two Skovgaard’s modifications (homoscedastic: 0.078 and 0.083; heteroschedastic: 0.089 and 0.093) and the Behrens-Fisher test (heteroscedastic: 0.179) claim that the average blood characteristics levels for male athletes competing in $T_{400}$ and $T_{Sprint}$ events are not significantly different at a 5% level. Instead, the $p$-values of the central limit theorem test (homoscedastic: 0.036) and the likelihood ratio test (homoscedastic: 0.027; heteroscedastic: 0.034) both lead to a different conclusion. Note that the central limit theorem test is not reliable with small $p$ and $n_{i}$, while the likelihood ratio test is not valid if $\kappa$ and $\tau$ are large. To assess the validity of the various testing procedures in this setting, we can also examine the empirical size of the various tests. We generate $R=10,000$ samples with sizes $n_{T_{400}}=18$ and $n_{T_{Sprint}}=11$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{T_{Sprint}}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{T_{Sprint}}\bar{y}_{T_{Sprint}})/(n_{T_{400}}+n_{T_{Sprint}})$ where $\bar{y}_{T_{400}}$ and $\bar{y}_{T_{Sprint}}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}^{-1}_{0}$, the pooled sample covariance matrix for the homoscedastic case, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$ sample covariance matrices of each group, for the heteroscedastic case. The simulation results for hypothesis (i) reported in Table 4 show that the directional test is more reliable than its competitors, having empirical size closer to the nominal level $0.05$. The central limit theorem test and likelihood ratio test suffer from large bias. These results confirm the simulation outputs in Section S3.
@@
Table 3: $p$-values of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
Table 4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
@@
For hypothesis (ii), all the tests used under heteroscedasticity in Table 3 give $p$-values directional: 0.097; likelihood ratio: 0.062; Skovgaard’s: 0.152 and 0.158) concluding that differences in mean blood characteristics among endurance-related events ($T_{400}$, $Row$ and $Swim$) for male athletes are not statistically significant at a 5% level. In the homoscedastic case, only the likelihood ratio test yields a significant $p$-value (0.045). Conversely, the other test statistics have $p$-values (directional: 0.092; central limit theorem: 0.069; Bartlett corrected: 0.084; Skovgaard’s: 0.101 and 0.105) that agree with the hypothesis of no significant difference among the three groups. Even in this case, a simulation study can be performed in order to verify the reliability of the considered methods. Under the null hypothesis, we generate $10,000$ samples with sizes $n_{T_{400}}=18$, $n_{Row}=15$ and $n_{Swim}=13$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{Row}\bar{y}_{Row}+n_{Swim}\bar{y}_{Swim})/(n_{T_{400}}+n_{Row}+n_{Swim})$ where $\bar{y}_{T_{400}}$, $\bar{y}_{Row}$ and $\bar{y}_{Swim}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}_{0}^{-1}$, the pooled sample covariance matrix under the homoscedasticity, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$, sample covariance matrices under the heteroscedasticity. Simulation results are reported in Table 4. In the homoscedastic case, the two Skovgaard’s modifications are conservative. The central limit theorem test and the likelihood test, on the other hand, have slightly inflated empirical size. Hence, the $p$-values of these four methods are not reliable in this scenario. In contrast, the directional test and the Bartlett corrected test well control the type I error. In the heteroscedastic case, the likelihood ratio test is overly inflated, whereas the remaining procedures, including the directional test, exhibit a good performance.